# Federated Class-Incremental Learning with Hierarchical Generative Prototypes

### v3

@@ 6 Conclusions
In this work, we propose Hierarchical Generative Prototypes (HGP), an approach aimed at mitigating Incremental and Federated Biases within Federated Class-Incremental Learning. HGP utilizes pre-trained architectures conditioned by prompting to achieve communication efficiency. We examine the bias implications of fine-tuning the entire model compared to using prompting techniques, demonstrating that prompting confines the bias primarily to the classification layer. Building on these insights, we propose Classifier Rebalancing as a solution. This method involves sampling features from a hierarchical Gaussian Mixture Model to train the classifier across all observed classes, effectively reducing bias in the classification layer. Through the integration of prompt learning and Classifier Rebalancing, we achieve superior performance compared to the State of The Art while transmitting only a minimal number of parameters.
@@
This work builds upon Prompt Learning, which effectively manages communication costs and Federated Bias while preserving adaptability. However, although leveraging a pre-trained model is a common strategy in Federated Learning literature (Guo et al. 2024; Liu et al. 2023; Bagwe et al. 2023), our findings may not extend to scenarios where models are trained from scratch. We leave the exploration of how other PEFT techniques impact federated bias to future research.
@@ Appendix A Communication cost
Efficient communication is paramount in Federated Learning because of the extensive coordination required between server and clients. In this section, we analyze the communication cost for each tested method. Figure A (left) illustrates the relationship between Final Average Accuracy (FAA) and communication cost for each method. The FAA results are derived from a quantity-based scenario with $\alpha=2$ on Tiny-ImageNet. The communication cost (reported in Figure A, right) is measured in Megabytes (MB) per communication round and represents the average data exchanged between a single client and the server ($C\rightarrow S$) and from the server to a single client ($S\rightarrow C$).
Notably, methods that do not utilize PEFT techniques incur significantly higher communication costs, as they optimize all parameters. Consequently, with larger models, communication costs escalate proportionally without corresponding performance gains. TARGET is the most expensive method overall because the server sends the entire model and the generated dataset to clients once per task. In contrast, the adoption of PEFT techniques enhances both efficiency and performance. L2P emerges as the most efficient method, using fewer prompts than its competitors. While CODA-P transmits slightly more parameters than L2P, it keeps communication costs comparable by only transmitting prompts for the current task and freezing the rest.
PLoRA is less efficient than prompt-based methods due to the larger size of LoRA modules. However, it still achieves costs that are two orders of magnitude lower than fine-tuning approaches while delivering superior performance. HGP demonstrates efficiency on par with the other prompt-based techniques, with slightly higher server-client communication cost. This is attributed to the Classifier Rebalancing process, which also changes the classification heads related to the previous tasks, thus necessitating to communicate the whole updated classifier back to the clients.
@@
Partition $\beta=0.5$ $\beta=0.1$ $\beta=0.05$ $\alpha=6$ $\alpha=4$ $\alpha=2$
Dataset CIFAR-100
L2PFL(BP) $77.4$ $(83.3)$ $71.3$ $(79.4)$ $69.4$ $(77.0)$ $72.8$ $(80.3)$ $70.4$ $(79.2)$ $63.4$ $(72.0)$
L2PFL $80.0$ $(83.7)$ $47.9$ $(57.7)$ $43.2$ $(52.1)$ $49.8$ $(60.1)$ $25.2$ $(41.9)$ $10.0$ $(24.3)$
Dataset Tiny-ImageNet
L2PFL(BP) $64.2$ $(66.9)$ $56.3$ $(52.5)$ $51.9$ $(43.2)$ $61.6$ $(58.0)$ $49.4$ $(39.3)$   $8.2$ $(10.2)$
L2PFL $60.3$ $(70.8)$ $42.2$ $(50.6)$ $36.5$ $(45.4)$ $51.4$ $(61.5)$ $26.8$ $(43.0)$   $7.5$ $(13.1)$
Table A: Batch-wise prompting. FAA []delimited-[][$\uparrow$][ ↑ ] and AIA []delimited-[][$\uparrow$][ ↑ ] for L2P, with and without batch-wise prompting at test time on all settings. AIA is reported between parenthesis.
Appendix B Implementation details
B.1 Hyperparameters
We utilize a pre-trained ViT-B/16 as the backbone for HGP and all the compared methods. Specifically, we initialize the models with supervised pre-trained weights on ImageNet-21K (Ridnik et al. 2021) for CIFAR-100 and self-supervised pre-trained weights from DINO (Caron et al. 2021) for Tiny-ImageNet. This selection is made to mitigate potential data leakage when adapting the model to the latter dataset, which is a subset of ImageNet (Deng et al. 2009).
In our method, we adopt a prompt pool composed of $10$ prompts for each task. Across all experiments, prompt components have a shape of $(8,d)$, where $d=768$ represents the embedding dimension for the chosen architecture. We apply these prompts using prefix tuning, conditioning the first $5$ layers of the backbone. For each task, we perform five communication rounds, during which clients observe their local datasets for five epochs with a batch size of $64$. Training is conducted using the Adam optimizer (Kingma and Ba 2015) with a learning rate of $0.001$; $\beta_{1}$ and $\beta_{2}$ are set to $0.9$ and $0.999$, respectively. A cosine annealing scheduler is employed to decay the learning rate during local training.
@@ B.1 Hyperparameters
Figure B: Real training image (left) and its reconstructions leveraging either the input image (center) or the generative prototype of the related class (right).
B.2 Metrics
We assess the performance of all methods using two widely adopted metrics in FCIL literature: Final Average Accuracy (FAA) and Average Incremental Accuracy (AIA).
@@ B.2 Metrics
On the other hand, AIA provides insight into the performance progression throughout the training process. It calculates the average accuracy after each task, considering all data encountered up to that point. After completion of the entire training procedure, the mean of all measurements is computed as follows:
$\text{AIA}=\frac{1}{T}\sum_{j=1}^{T}\left[\frac{1}{j}\sum_{i=1}^{j}A_{i}^{j}\right].$
Appendix C Batch-wise prompting
To maintain consistency with other studies, in the main paper we report L2P performance using Batch-wise Prompting (BP) at test time (Jung et al. 2023; Smith et al. 2023). Such a procedure involves selecting the same set of prompts for all samples within a single batch during testing, thus violating the test assumption of processing each sample independently. This gives L2P an unfair advantage compared to other methods when the test dataset is not shuffled: i.e., when the ground-truth labels of the examples within a single batch are typically the same.
In this respect, Table A presents L2P’s performance with and without batch-wise prompting. Without BP, the set of prompts is selected individually for each sample, ensuring independent predictions. L2P achieves similar results with and without batch-wise prompting when $\beta=0.5$. However, in all the other settings, batch-wise prompting appears to increment L2P’s performance considerably, especially on CIFAR-100.
@@ Appendix C Batch-wise prompting
Appendix D Protypes of features are privacy-preserving
To verify the safety of generative prototypes, we firstly examine a hypothetical scenario where an attacker gains access to the trained model along with a real image utilized during training. In such a case, the attacker could potentially employ the feature inversion technique outlined in (Zhao et al. 2020). This methodology provides for training a autoencoder to precisely match the input image. Specifically, the autoencoder takes fixed noise as input and reconstructs the image. Then, both the real and reconstructed image are passed through the frozen backbone. The objective is to minimize the mean squared error (MSE) loss between their respective features. Figure B (left) shows the real training image from the Tiny-ImageNet dataset, while Figure B (center) displays the reconstruction obtained by the trained autoencoder. Through this process, no significant semantics can be recovered. At most, the class might be guessed, but the image does not provide any detail of the original sample.
In another setting, we suppose that the attacker gains access to the centralized server, which has visibility into clients’ generative prototypes. To recover an input image, initial random noise in the input space is optimized by minimizing MSE between its features and those of a fixed sample from the generative prototype. Figure B (right) shows that this results in a shapeless, noisy reconstruction, making it impracticable for an attacker to recreate real images from these feature distributions.
Appendix E Algorithm

### v4

@@ 6 Conclusions
In this work, we propose Hierarchical Generative Prototypes (HGP), an approach aimed at mitigating Incremental and Federated Biases within Federated Class-Incremental Learning. HGP utilizes pre-trained architectures conditioned by prompting to achieve state-of-the-art performance while maintaining parameter efficiency. To tackle the problem of Federated Bias and Incremental Bias, we examine the implications of fine-tuning the entire model compared to using prompting techniques, demonstrating that the latter confines bias to the classification layer. Building on these insights, we propose Classifier Rebalancing – i.e., sampling features from a hierarchical Gaussian Mixture Model to train the classifier across all observed classes – as an effective solution. Through the integration of prompt learning and Classifier Rebalancing, we achieve SOTA performance while learning only a minimal number of parameters.
@@
This work builds upon Prompt Learning, which allows adapting a pre-trained architecture on a downstream task. Although leveraging a pre-trained model is a common strategy in Federated Learning literature guo2024federated ; liu2023fedet ; bagwe2023fed , our findings may not extend to scenarios where models are trained from scratch. Moreover, although prompt tuning effectively localizes bias to the classifier layer, the behavior of other PEFT techniques – such as LoRA or Adapters – under continual decentralized scenarios remains unexplored. Investigating these methods may uncover important trade-offs between efficiency and bias control, which we leave for future work. Finally, assessing HGP’s robustness to adversarial settings with malicious clients remains an important direction, as it is critical for ensuring reliable deployment in real-world applications.
@@ Appendix A Communication Cost
Efficient communication is paramount in Federated Learning because of the extensive coordination required between server and clients. In this section, we analyze illustrates the relationship between Final Average Accuracy (FAA) and communication cost for each method. The FAA results are provided for CUB-200 (Figure B), Imagenet-R, and CIFAR-100 (Figure A), with $\beta=0.05$, and communication costs are measured in Megabytes (MB) per communication round, representing the average data exchanged between a single client and the server.
Notably, methods that do not utilize PEFT techniques incur significantly higher communication costs, as they optimize all parameters. Consequently, with larger models, communication costs escalate proportionally without corresponding performance gains. TARGET is the most expensive method overall because the server sends the entire model and the generated dataset to clients once per task. In contrast, the adoption of PEFT techniques enhances both efficiency and performance. L2P emerges as the most efficient method, using fewer prompts than its competitors, followed by DualPrompt and CODA-P. PILoRA is on par to the previous prompt-based solutions in terms of efficiency. HGP demonstrates efficiency comparable with the PEFT techniques, ensuring a strong performance gain.
@@
Appendix B Computational Complexity
Our experimental setup aligns with prior work, assuming clients can fine-tune transformer-based networks. However, our method reduces training cost significantly by using prompt tuning, updating only $\sim$1.5of the total model parameters.
To ensure efficiency and numerical stability, we employ diagonal covariance matrices (represented as vectors) as prototypes. Empirically, full covariance matrices provide no accuracy gains and incur higher computation and memory costs.
[…]
In total, the sampling cost is approximately $770$ FLOPs for ViT-B/16. For comparison, a single forward pass on ViT-B/16 requires $\sim$17 GFLOPs, making our sampling overhead negligible. Moreover, inference time is further reduced by our proposed prompting technique, which employs a single shared prompt rather than a prompt pool thus decreasing inference cost by up to 50%.
Appendix C Implementation Details
C.1 Hyperparameters
We utilize a pre-trained ViT-B/16 as the backbone for HGP and all the compared methods. Specifically, we initialize the models with supervised pre-trained weights on ImageNet-21K [41] for all datasets.
In our method, we adopt a single prompt for all tasks, negating both the need for a double forward pass while also avoiding the need to store old task-specific parameters . Across all experiments, prompt components have a shape of $(200,d)$, where $d=768$ represents the embedding dimension for the chosen architecture. We train said prompt using prefix tuning, conditioning the first $5$ layers of the backbone. For each task, we perform five communication rounds, during which clients observe their local datasets for five epochs with a batch size of $16$. Training is conducted using the Adam optimizer [20] with a learning rate of $0.003$; $\beta_{1}$ and $\beta_{2}$ are set to $0.9$ and $0.999$, respectively.
@@ C.1 Hyperparameters
The specific training details for all implemented methods are documented in Appendix H.
C.2 Metrics
We assess the performance of all methods using the most commonly used metric in FCIL literature: Final Average Accuracy (FAA).
@@ C.2 Metrics
Figure C: Real training image (left) and its reconstructions leveraging either the input image (center) or the generative prototype of the related class (right).
Appendix D Protypes of Features are Privacy-preserving
To verify the safety of generative prototypes, we firstly examine a hypothetical scenario where an attacker gains access to the trained model along with a real image utilized during training. In such a case, the attacker could potentially employ the feature inversion technique outlined in [58]. This methodology provides for training a autoencoder to precisely match the input image. Specifically, the autoencoder takes fixed noise as input and reconstructs the image. Then, both the real and reconstructed image are passed through the frozen backbone. The objective is to minimize the mean squared error (MSE) loss between their respective features. Figure C (left) shows the real training image from the Tiny-ImageNet dataset, while Figure C (center) displays the reconstruction obtained by the trained autoencoder. Through this process, no significant semantics can be recovered. At most, the class might be guessed, but the image does not provide any detail of the original sample.
In another setting, we suppose that the attacker gains access to the centralized server, which has visibility into clients’ generative prototypes. To recover an input image, initial random noise in the input space is optimized by minimizing MSE between its features and those of a fixed sample from the generative prototype. Figure C (right) shows that this results in a shapeless, noisy reconstruction, making it impracticable for an attacker to recreate real images from these feature distributions.
Appendix E Scalability to Large Client Populations
To assess the scalability of our approach, we conducted additional experiments on CIFAR-100 using $100$ clients with varying participation rates under extreme distribution imbalance ($\beta=0.05$). Specifically, we tested participation rates of $10\%$, $20\%$, $50\%$, and $100\%$, obtaining average accuracies of $80.82\%$, $87.33\%$, $89.51\%$, and $90.08\%$, respectively. These results demonstrate that our method remains robust even under the most challenging federated settings, with only $10$ clients participating per round. Remarkably, in this low-participation regime, our method still outperforms all competing methods evaluated with just $10$ clients and full participation.
Appendix F Algorithm
@@
Appendix G Standard Deviations
The standard deviations across three runs for each method are presented in Table A and Table B.
Table A: Standard deviations (reported in FAA) for CIFAR-100, ImageNet-R and ImageNet-A.
CIFAR-100 ImageNet-R ImageNet-A
Distrib. $\boldsymbol{\beta}$ $0.5$ $0.1$ $0.05$ $0.5$ $0.1$ $0.05$ $1.0$ $0.5$ $0.2$
EWC $1.38$ $2.32$ $2.67$ $1.69$ $1.14$ $1.14$ $1.34$ $0.75$ $1.38$
LwF $1.15$ $2.38$ $4.25$ $1.91$ $0.93$ $1.38$ $0.30$ $0.27$ $1.61$
FisherAVG $0.40$ $3.67$ $3.18$ $0.77$ $0.47$ $0.93$ $1.59$ $1.36$ $1.49$
RegMean $0.01$ $1.78$ $4.87$ $1.10$ $0.85$ $1.44$ $0.59$ $1.10$ $0.70$
CCVR $1.13$ $2.49$ $1.39$ $0.57$ $0.49$ $2.29$ $1.19$ $0.63$ $2.26$
L2P $1.62$ $1.49$ $1.85$ $1.17$ $1.61$ $1.35$ $0.72$ $0.81$ $2.01$
DualPrompt $2.17$ $3.33$ $1.42$ $0.83$ $1.68$ $1.35$ $0.34$ $1.70$ $0.86$
CODA-P $1.15$ $1.96$ $0.96$ $0.98$ $1.95$ $1.19$ $0.16$ $1.63$ $0.22$
FedProto $2.05$ $0.94$ $1.96$ $2.09$ $2.28$ $0.72$ $1.18$ $0.71$ $2.17$
TARGET $0.79$ $1.02$ $1.81$ $0.33$ $1.58$ $1.21$ $0.65$ $0.46$ $1.34$
PILoRA $0.28$ $0.98$ $4.08$ $0.29$ $1.01$ $0.33$ $0.08$ $0.31$ $0.33$
PIP $0.29$ $1.32$ $0.68$ $0.52$ $0.48$ $0.34$ $0.78$ $0.20$ $0.83$
LoRM $0.27$ $0.20$ $0.79$ $0.04$ $0.16$ $0.87$ $0.86$ $0.75$ $0.68$
HGP $0.08$ $0.14$ $0.24$ $0.43$ $0.39$ $0.31$ $0.90$ $0.91$ $0.29$
Table B: Standard deviations (reported in FAA) for EuroSAT, Cars-196, and CUB-200.
EuroSAT Cars-196 CUB-200
Distrib. $\boldsymbol{\beta}$ $1.0$ $0.5$ $0.2$ $1.0$ $0.5$ $0.2$ $1.0$ $0.5$ $0.2$
EWC $7.33$ $5.78$ $6.60$ $1.72$ $0.46$ $1.00$ $0.55$ $0.94$ $1.68$
LwF $3.32$ $4.58$ $5.12$ $1.35$ $2.81$ $2.01$ $2.07$ $2.15$ $2.02$
FisherAVG $5.43$ $6.07$ $3.40$ $2.10$ $1.78$ $0.78$ $0.26$ $1.35$ $2.21$
RegMean $3.99$ $7.21$ $5.68$ $0.23$ $1.11$ $1.80$ $1.79$ $2.63$ $2.27$
CCVR $9.01$ $7.16$ $6.26$ $1.49$ $0.87$ $2.08$ $1.22$ $1.69$ $1.73$
L2P $2.52$ $3.49$ $2.02$ $0.87$ $1.94$ $0.36$ $2.38$ $0.78$ $1.21$
DualPrompt $8.75$ $6.70$ $8.89$ $0.87$ $1.79$ $1.82$ $0.42$ $2.62$ $2.69$
CODA-P $4.52$ $6.27$ $6.54$ $1.05$ $2.24$ $1.89$ $0.51$ $1.52$ $1.86$
FedProto $3.58$ $8.74$ $7.96$ $1.11$ $0.35$ $0.87$ $0.96$ $0.67$ $1.99$
TARGET $4.12$ $6.20$ $5.31$ $0.93$ $0.68$ $1.54$ $1.17$ $0.65$ $2.06$
PILoRA $4.45$ $4.07$ $4.27$ $0.21$ $0.33$ $0.18$ $0.51$ $0.28$ $0.51$
PIP $5.95$ $0.49$ $2.61$ $1.11$ $2.15$ $4.01$ $1.90$ $1.10$ $1.42$
LoRM $1.75$ $3.34$ $6.51$ $1.07$ $0.26$ $0.92$ $0.86$ $0.98$ $0.44$
HGP $1.03$ $1.20$ $1.37$ $1.73$ $1.58$ $0.30$ $0.36$ $0.19$ $0.60$
Appendix H Hyperparameter Tables
The following acronyms and symbols are used throughout the paper: lr denotes the learning rate; $\textit{lr}_{pr}$ refers to the learning rate applied to prototypes; $\lambda_{\text{KL}}$ is the weighting factor for the Knowledge Distillation (KL) loss; r represents the rank used in low-rank matrix approximations; $\textit{g}_{ep}$ indicates the number of epochs used for training and sample generation in the generator network; and $\gamma$ is the decay coefficient applied to the off-diagonal elements of the Gram matrices for the backbone (first component) and classifier (second component).
H.1 CIFAR-100
$\boldsymbol{\beta}$ $0.5$ $0.1$ $0.05$
EwC lr: 1e-5 lr: 1e-5 lr: 1e-5
LwF lr: 1e-5 lr: 1e-5 lr: 1e-5
FisherAVG lr: 1e-5 lr: 1e-5 lr: 1e-5
RegMean lr: 1e-5; $\gamma$: (0.5, 0.5) lr: 1e-5; $\gamma$: (0.5, 0.5) lr: 1e-5; $\gamma$: (0.5, 0.5)
CCVR lr: 1e-5 lr: 1e-5 lr: 1e-5
L2P lr: 3e-2 lr: 3e-2 lr: 3e-2
DualPrompt lr: 5e-2 lr: 5e-2 lr: 5e-2
CODA-P lr: 1e-3 lr: 1e-3 lr: 1e-3
FedProto lr: 1e-5 lr: 1e-5 lr: 1e-5
TARGET lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30
PILoRA lr: 2e-2; $lr_{pr}$: 1e-4 lr: 2e-2; $lr_{pr}$: 1e-4 lr: 2e-2; $lr_{pr}$: 1e-4
PIP lr: 5e-3 lr: 5e-3 lr: 5e-3
LoRM lr: 3e-4; r: 1 $\gamma$: (0, 0.5) lr: 1e-4; r: 16 $\gamma$: (0, 0.5) lr: 5e-4; r: 16 $\gamma$: (0, 0.5)
HGP lr: 3e-3 lr: 3e-3 lr: 3e-3
H.2 ImageNet-R
$\boldsymbol{\beta}$ $0.5$ $0.1$ $0.05$
EwC lr: 1e-5 lr: 1e-5 lr: 1e-5
LwF lr: 1e-5 lr: 1e-5 lr: 3e-5
FisherAVG lr: 1e-5 lr: 1e-5 lr: 1e-5
RegMean lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1)
CCVR lr: 1e-5 lr: 1e-5 lr: 1e-5
L2P lr: 3e-2 lr: 3e-2 lr: 3e-2
DualPrompt lr: 3e-4 lr: 3e-4 lr: 3e-4
CODA-P lr: 1e-3 lr: 1e-3 lr: 1e-3
FedProto lr: 1e-5 lr: 1e-5 lr: 3e-5
TARGET lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30
PILoRA lr: 2e-2; $\textit{lr}_{pr}$: 1e-4 lr: 2e-2; $\textit{lr}_{pr}$: 1e-4 lr: 2e-2; $\textit{lr}_{pr}$: 1e-4
PIP lr: 3e-4 lr: 3e-4 lr: 3e-4
LoRM lr: 3e-3; r: 2 $\gamma$: (0, 0.5) lr: 1e-3; r: 32 $\gamma$: (0, 0.5) lr: 1e-3; r: 16 $\gamma$: (0, 0.5)
HGP lr: 3e-2 lr: 3e-2 lr: 1e-2
H.3 ImageNet-A
$\boldsymbol{\beta}$ $1.0$ $0.5$ $0.2$
EwC lr: 1e-5 lr: 1e-5 lr: 1e-5
LwF lr: 1e-5 lr: 1e-5 lr: 3e-5
FisherAVG lr: 1e-5 lr: 1e-5 lr: 1e-5
RegMean lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1)
CCVR lr: 1e-5 lr: 1e-5 lr: 1e-5
L2P lr: 3e-2 lr: 3e-2 lr: 3e-1
DualPrompt lr: 3e-2 lr: 3e-2 lr: 3e-2
CODA-P lr: 1e-2 lr: 1e-2 lr: 1e-2
FedProto lr: 3e-5 lr: 1e-5 lr: 1e-5
TARGET lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30
PILoRA lr: 2e-2; $\textit{lr}_{pr}$ 1e-4 lr: 1e-2; $\textit{lr}_{pr}$: 1e-4 lr: 2e-2; $\textit{lr}_{pr}$: 1e-4
PIP lr: 3e-2 lr: 3e-2 lr: 3e-2
LoRM lr: 1e-2; r: 4 $\gamma$: (0, 0.5) lr: 1e-2; r: 4 $\gamma$: (0, 0.5) lr: 1e-2; r: 4 $\gamma$: (0, 0.5)
HGP lr: 3e-2 lr: 3e-2 lr: 3e-2
H.4 EuroSAT
$\boldsymbol{\beta}$ $1.0$ $0.5$ $0.2$
EwC lr: 1e-5 lr: 1e-5 lr: 1e-5
LwF lr: 1e-5 lr: 1e-5 lr: 3e-5
FisherAVG lr: 1e-5 lr: 1e-5 lr: 1e-5
RegMean lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1)
CCVR lr: 1e-5 lr: 1e-5 lr: 1e-5
L2P lr: 3e-2 lr: 3e-2 lr: 3e-2
DualPrompt lr: 1e-3 lr: 1e-3 lr: 1e-3
CODA-P lr: 1e-3 lr: 1e-3 lr: 1e-3
FedProto lr: 3e-5 lr: 1e-5 lr: 1e-5
TARGET lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-5; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30
PILoRA lr: 2e-2; $\textit{lr}_{pr}$: 1e-4 lr: 2e-2; $\textit{lr}_{pr}$: 1e-4 lr: 2e-2; $\textit{lr}_{pr}$: 1e-4
PIP lr: 1e-3 lr: 1e-3 lr: 1e-3
LoRM lr: 3e-3; r: 1 $\gamma$: (0, 0.5) lr: 3e-3; r: 1 $\gamma$: (0, 0.5) lr: 1e-3; r: 4 $\gamma$: (0, 0.5)
HGP lr: 1e-3 lr: 3e-4 lr: 1e-4
H.5 Cars-196
$\boldsymbol{\beta}$ $1.0$ $0.5$ $0.2$
EwC lr: 1e-5 lr: 1e-5 lr: 1e-5
LwF lr: 1e-5 lr: 1e-5 lr: 3e-5
FisherAVG lr: 1e-5 lr: 1e-5 lr: 1e-5
RegMean lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1)
CCVR lr: 1e-5 lr: 1e-5 lr: 1e-5
L2P lr: 3e-2 lr: 3e-2 lr: 3e-2
DualPrompt lr: 1e-1 lr: 1e-1 lr: 1e-1
CODA-P lr: 3e-2 lr: 3e-2 lr: 3e-2
FedProto lr: 1e-5 lr: 1e-5 lr: 1e-5
TARGET lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30
PILoRA lr: 1e-1; $\textit{lr}_{pr}$ 1e-4 lr: 1e-1; $\textit{lr}_{pr}$: 1e-4 lr: 1e-1; $\textit{lr}_{pr}$: 1e-4
PIP lr: 1e-1 lr: 1e-1 lr: 1e-1
LoRM lr: 1e-2; r: 8 $\gamma$: (0, 0.5) lr: 1e-2; r: 8 $\gamma$: (0, 0.5) lr: 1e-2; r: 4 $\gamma$: (0, 0.5)
HGP lr: 3e-3 lr: 3e-3 lr: 3e-3
H.6 CUB-200
$\boldsymbol{\beta}$ $1.0$ $0.5$ $0.2$
EwC lr: 1e-5 lr: 1e-5 lr: 1e-5
LwF lr: 1e-5 lr: 1e-5 lr: 3e-5
FisherAVG lr: 1e-5 lr: 1e-5 lr: 1e-5
RegMean lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1) lr: 1e-5; $\gamma$: (0.1, 0.1)
CCVR lr: 1e-5 lr: 1e-5 lr: 1e-5
L2P lr: 3e-1 lr: 3e-1 lr: 3e-1
DualPrompt lr: 1e-1 lr: 1e-1 lr: 1e-1
CODA-P lr: 1e-3 lr: 1e-3 lr: 1e-3
FedProto lr: 1e-5 lr: 1e-5 lr: 1e-5
TARGET lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30 lr: 1e-4; $\lambda_{\text{KL}}$: 25; $\textit{g}_{ep}$: 30
PILoRA lr: 1; $\textit{lr}_{pr}$ 1e-4 lr: 1; $\textit{lr}_{pr}$: 1e-4 lr: 1; $\textit{lr}_{pr}$: 1e-4
PIP lr: 1e-1 lr: 1e-1 lr: 1e-1
LoRM lr: 1e-2; r: 1 $\gamma$: (0, 0.3) lr: 3e-2; r: 1 $\gamma$: (0, 0.3) lr: 3e-2; r: 1 $\gamma$: (0, 0.3)
HGP lr: 1e-1 lr: 1e-1 lr: 1e-1